# Tracially lyriform $\mathrm{C}^{*}$-algebras

### v1

@@ Introduction
This article is also inspired by spiders. As Barth writes in [Bar02, Chapter VIII], “The world in which spiders live is a world full of vibrations, and their vibration sense is correspondingly well developed.” These vibrations in the world of an arachnid are detected via slit sense organs embedded in their exoskeletons. In contrast to other arachnids like scorpions, whip spiders and whip scorpions (see [Bar02, p.41]), spiders are unique in the tendency of slit organs near their leg joints to be clustered into closely spaced parallel arrangements called lyriform organs. One such arrangement is displayed in Figure 1, which shows a female cross orb-weaver (Araneus diadematus) and a selection of her anatomy (cf. [MTM22, Fig. 1]). The picture explains the terminology: ‘lyriform’ means ‘lyre shaped’. The different lyriform organs distributed across the spider’s legs play exteroceptive or proprioceptive roles, that is, they give information about the spider’s environment and her positioning within it. That is also the purpose of lyriform structure in the $\mathrm{C}^{*}$-world. An inductive limit of Wasserstein spheres has the function of a tracial exoskeleton, with the lyriform structure providing geometric information about the way that the underlying $\mathrm{C}^{*}$-algebra is embedded in its surroundings.
@@ Acknowledgements
  • $\cdot$
    If $A$ and $B$ are unital $\mathrm{C}^{*}$-algebras, then $\operatorname{\mathrm{Emb}}(A,B)$ denotes the set of unital embeddings (that is, injective -homomorphisms) $A\to B$. Two -homomorphisms $\varphi,\psi\colon A\to B$ are said to be approximately unitarily equivalent, written $\varphi\sim_{au}\psi$, if for every $\varepsilon>0$ and every finite set $F\subseteq A$ there is a unitary $u\in B$ such that $\|\varphi(a)-u\psi(a)u^{*}\|<\varepsilon$ for every $a\in F$. If $A$ is separable, this is equivalent to the existence of a sequence $(u_{n})_{n\in\mathbb{N}}$ of unitaries in $B$ such that $\varphi(a)=\lim_{n\to\infty}u_{n}\psi(a)u_{n}^{*}$ for every $a\in A$.
  • $\cdot$
@@ 2. Tracially lyriform $\mathrm{C}^{*}$-algebras
The underlying $\mathrm{C}^{*}$-algebra is the quantum permutation group $C(S_{n}^{+})$ introduced in [Wan98]. Anshu, Jekel and Landry equip the trace space $T(C(S_{n}^{+}))$ with metrics called free Hamming distances, based on the Biane–Voiculescu notion of a tracial coupling. Such a coupling between traces $\tau_{1}$ and $\tau_{2}$ on a $\mathrm{C}^{*}$-algebra $A$ means a tuple $(M,\tau,\alpha_{1},\alpha_{2})$, where $(M,\tau)$ is a tracial von Neumann algebra and, for $i=1,2$, $\alpha_{i}\colon A\to M$ is a -homomorphism with $\tau\circ\alpha_{i}=\tau_{i}$. The two versions of free Hamming distance $W_{H,\wedge}(\tau_{1},\tau_{2})$ and $W_{H,1}(\tau_{1},\tau_{2})$ are defined in [AJL25, Definitions 1 and 2] as infima over all tracial couplings of expressions involving $(M,\tau,\alpha_{1},\alpha_{2})$ and the canonical generating projections $\{u_{i,j}\}_{1\leq i,j\leq n}$ of $C(S_{n}^{+})$, namely:
@@ 4. Inductive limits and the Lyre category
    $(\rho_{n})_{n\in\mathbb{N}}$ is uniformly effectively quasiconvex, meaning that, in the definition (2.3) of effective quasiconvexity, a single function $h$ can be used for all of the metrics $\rho_{n}$, and
@@
$\rho_{\infty}((\sigma_{n})_{n=1}^{\infty},(\tau_{n})_{n=1}^{\infty})=\sup_{n\in\mathbb{N}}\rho_{n}(\sigma_{n},\tau_{n}).$ (4.1)
@@
The listed hypotheses guarantee that $\rho_{\infty}$ is effectively quasiconvex and that $\operatorname{\mathrm{diam}}(T(A),\rho_{\infty})<\infty$. They hold for tracial Prokhorov spaces and tracial Wasserstein spaces of bounded diameter by Proposition 3.1. The expression (4.1) is lower semicontinuous, as it defines $\rho_{\infty}$ as a pointwise supremum of an increasing sequence of lower semicontinuous functions on $T(A)$. It also ensures that the inclusion maps $\varphi_{n,\infty}\colon A_{n}\to A$ are tracially nonexpansive. In particular, $\rho_{n}$-Lipschitz elements in any $A_{n}$ are mapped to $\rho_{\infty}$-Lipschitz elements in the limit, so $L_{\rho_{\infty}}$ is indeed densely finite. This verifies that $(A,\rho_{\infty})$ is a tracially lyriform $\mathrm{C}^{*}$-algebra. Suppose next that $(B,\rho_{B})$ is tracially lyriform and $(\psi_{n}\colon A_{n}\to B)_{n\in\mathbb{N}}$ is a compatible system of $\mathbf{Lyre}$-morphisms. Let $\psi\colon A\to B$ be the induced unital -monomorphism. Then, for $\sigma,\tau\in T(B)$ăwe have
@@ 6.1. Noncommutative continuous transport
The continuous transport constant of a tracial Wasserstein space $(A,W_{\mathfrak{p},\rho})$ observing a compact metric space $(X,\rho)$ is the minimal element $k=k_{(A,\rho)}$ of $1,\infty]$ for which there exists a nucleus $\mathcal{L}(A)$ with the following property: For every pair of faithful traces $\sigma,\tau\in T(A)$ and every $\varepsilon>0$, there is an automorphism $\alpha$ of $A$ that is homotopic to the identity and satisfies
@@
This in particular implies that the ‘noncommutative’ transport constant $k_{C(X)}$ of $(C(X),\rho)$ (in the sense of Definition 6.1) agrees with the ‘commutative’ transport constant $k_{X}$ introduced in [Jac23b, Definition 2.6]. Actually, that is not quite true: Definition 6.1 is stated for faithful measures while [Jac23b, Definition 2.6] only considers measures that are good, that is, faithful as well as diffuse. So, formally, we have $k_{C(X)}\geq k_{X}$. The difference really becomes substantive when one is interested in an exact transport map rather than an approximate one, and for our present needs is insignificant. Indeed, as observed in [Jac23b, Remark 2.7], our typical method of transport does not depend on diffuseness (see the proof of Theorem 6.4), so there is no practical difference between the two notions. Moreover, as good measures are generic in $\operatorname{\mathcal{M}}(X)$, faithful measures can be perturbed to good ones (see the proof of [Jac23b, Theorem 3.1]). A final point in favour of working with faithful traces, as discussed at the beginning of this section, is that these are what actually parameterise embeddings.
@@
Let $X$ be a compact, connected Riemannian manifold of dimension at least three and let $\rho$ be an intrinsic metric on $X$ that is Lipschitz equivalent to the intrinsic metric induced by the Riemannian structure. Let $X_{p,q}$ be a generalised dimension drop algebra over $X$ for some $p,q\in\mathbb{N}$ and $x_{0}\neq x_{1}\in X$. Then, $k_{(X_{p,q},\rho)}=1$.
@@ 6.2. Distances between unitary orbits
$d_{U}(\varphi,\psi)|_{\mathcal{L}(A)}:=\adjustlimits{\inf}_{u\in\mathcal{U}(B)}{\sup}_{F\subset\joinrel\subset\mathcal{L}(A)}\sup_{a\in F}\|\varphi(a)-u\psi(a)u^{*}\|.$ (6.10)
@@
The inequality (6.12) is proved in exactly the same way as [JSV21, Theorem 4.11] and its generalisation [Jac23b, Theorem 3.1], which cover commutative domains $A=C(X)$. For (6.13), since we know from Theorem 6.4 that the canonical nucleus $\mathcal{D}_{r}(A)$ defined in (6.4) witnesses $k_{(A,\rho)}=1$, it then suffices to prove that $d_{U}(\varphi,\psi)|_{\mathcal{D}_{r}(A)}\geq W_{\infty}(\varphi,\psi)$. This can also be deduced from [JSV21, Jac23b], since the embedding $\iota$ of $C(X)=X_{1,1}$ into the centre of $A=X_{p,q}$ restricts to an inclusion of canonical nuclei. It follows that
@@
$\overrightarrow{W}_{\infty,\omega}((\sigma_{k})_{k=1}^{\infty},(\tau_{k})_{k=1}^{\infty})=\sup_{k\in\mathbb{N}}\overrightarrow{W}_{\infty,\omega_{k}}(\sigma_{k},\tau_{k})$ (6.17)
@@
$\overrightarrow{W}_{\infty,\rho}((\sigma_{k})_{k=1}^{\infty},(\tau_{k})_{k=1}^{\infty})=\sup_{k\in\mathbb{N}}\overrightarrow{W}_{\infty,\rho_{k}}(\sigma_{k},\tau_{k}).$ (6.18)

### v2

@@ Abstract.
Key words and phrases:
Quantum metric spaces, $\mathrm{C}^{*}$-classification, Alexandrov spaces, metric geometry, optimal transport.
2020 Mathematics Subject Classification:
Primary 46L35; Secondary 53C23, 28A33, 49Q22
@@ Introduction
This article is also inspired by spiders. As Barth writes in [Bar02, Chapter VIII], “The world in which spiders live is a world full of vibrations, and their vibration sense is correspondingly well developed.” Arachnids detect these vibrations via slit sense organs embedded in their exoskeletons. In contrast to others in the class Arachnida, like scorpions, whip spiders and whip scorpions (see [Bar02, p.41]), spiders are unique in the tendency of slit organs near their leg joints to be clustered into closely spaced parallel arrangements called lyriform organs. One such arrangement is displayed in Figure 1, which shows a female cross orb-weaver (Araneus diadematus) and a selection of her anatomy (cf. [MTM22, Fig. 1]). The picture explains the terminology: ‘lyriform’ means ‘lyre shaped’. The different lyriform organs distributed across the spider’s legs play exteroceptive or proprioceptive roles, that is, they give information about the spider’s environment and her positioning within it. That is also the purpose of lyriform structure in the $\mathrm{C}^{*}$-world. An inductive limit of Wasserstein spheres has the function of a tracial exoskeleton, with the lyriform structure providing geometric information about the way that the underlying $\mathrm{C}^{*}$-algebra is embedded in its surroundings.
@@ Acknowledgements
[…]
    If $A$ and $B$ are unital $\mathrm{C}^{*}$-algebras, then $\operatorname{\mathrm{Emb}}(A,B)$ denotes the set of unital embeddings (that is, injective -homomorphisms) $A\to B$. Two -homomorphisms $\varphi,\psi\colon A\to B$ are said to be approximately unitarily equivalent, written $\varphi\sim_{au}\psi$, if for every $\varepsilon>0$ and every finite set $F\subseteq A$ there is a unitary $u\in B$ such that $\|\varphi(a)-u\psi(a)u^{*}\|<\varepsilon$ for every $a\in F$. If $A$ is separable, this is equivalent to the existence of a sequence $(u_{n})_{n\in\mathbb{N}}$ of unitaries in $B$ such that $\varphi(a)=\lim_{n\to\infty}u_{n}\psi(a)u_{n}^{*}$ for every $a\in A$. We frequently rely on both the existence [CGSTW23, Corollary C and Theorem 9.9] and uniqueness [CGSTW23, Theorem B] aspects of classification to lift compatible morphisms between ordered $K$-theory $(K_{0}(\cdot),K_{0}(\cdot)_{+},[1_{(\cdot)}],K_{1}(\cdot))$ and traces $T(\cdot)$ to unital embeddings, uniquely up to approximate unitary equivalence in situations where the total invariant $\underline{K}T_{u}$ developed in [CGSTW23] reduces to ordered $K$-theory and traces.
[…]
@@ 2. Tracially lyriform $\mathrm{C}^{*}$-algebras
The underlying $\mathrm{C}^{*}$-algebra is the quantum permutation group $C(S_{n}^{+})$ introduced in [Wan98]. Anshu, Jekel and Landry equip the trace space $T(C(S_{n}^{+}))$ with metrics called free Hamming distances, based on the Biane–Voiculescu notion of a tracial coupling. Such a coupling between traces $\tau_{1}$ and $\tau_{2}$ on a $\mathrm{C}^{*}$-algebra $A$ means a tuple $(M,\tau,\alpha_{1},\alpha_{2})$, where $(M,\tau)$ is a tracial von Neumann algebra and, for $i=1,2$, $\alpha_{i}\colon A\to M$ is a -homomorphism with $\tau\circ\alpha_{i}=\tau_{i}$. The two versions of free Hamming distance $W_{H,\wedge}(\tau_{1},\tau_{2})$ and $W_{H,1}(\tau_{1},\tau_{2})$ are defined in [AJL25, Definitions 1.2 and 1.3] as infima over all tracial couplings of expressions involving $(M,\tau,\alpha_{1},\alpha_{2})$ and the canonical generating projections $\{u_{i,j}\}_{1\leq i,j\leq n}$ of $C(S_{n}^{+})$, namely:
@@ 4. Inductive limits and the Lyre category
    $\sup_{n\in\mathbb{N}}\operatorname{\mathrm{diam}}(T(A_{n}),\rho_{n})<\infty$, and
@@
    $\rho_{\infty}((\sigma_{n})_{n=1}^{\infty},(\tau_{n})_{n=1}^{\infty})=\sup_{n\in\mathbb{N}}\rho_{n}(\sigma_{n},\tau_{n})$ (4.1)
    is effectively quasiconvex (which holds whenever $(\rho_{n})_{n\in\mathbb{N}}$ is uniformly effectively quasiconvex, meaning that, in the definition (2.3) of effective quasiconvexity, a single function $h$ can be used for all of the metrics $\rho_{n}$).
@@
The listed hypotheses guarantee that $\rho_{\infty}$ is effectively quasiconvex and $\operatorname{\mathrm{diam}}(T(A),\rho_{\infty})<\infty$. By Proposition 3.1, both conditions hold for tracial Prokhorov spaces and for tracial $\mathfrak{p}$-Wasserstein spaces of bounded diameter. The expression (4.1) is lower semicontinuous, as it defines $\rho_{\infty}$ as a pointwise supremum of an increasing sequence of lower semicontinuous functions on $T(A)$. It also ensures that the inclusion maps $\varphi_{n,\infty}\colon A_{n}\to A$ are tracially nonexpansive. In particular, $\rho_{n}$-Lipschitz elements in any $A_{n}$ are mapped to $\rho_{\infty}$-Lipschitz elements in the limit, so $L_{\rho_{\infty}}$ is indeed densely finite. This verifies that $(A,\rho_{\infty})$ is a tracially lyriform $\mathrm{C}^{*}$-algebra. Suppose next that $(B,\rho_{B})$ is tracially lyriform and $(\psi_{n}\colon A_{n}\to B)_{n\in\mathbb{N}}$ is a compatible system of $\mathbf{Lyre}$-morphisms. Let $\psi\colon A\to B$ be the induced unital -monomorphism. Then, for $\sigma,\tau\in T(B)$ăwe have
@@ 6.1. Noncommutative continuous transport
The continuous transport constant of a tracial Wasserstein space $(A,W_{\mathfrak{p},\rho})$ relative to a nucleus $\mathcal{L}(A)$ is the minimal element $k=k_{(A,\rho)}$ of $1,\infty]$ with the following property: For every pair of faithful traces $\sigma,\tau\in T(A)$ and every $\varepsilon>0$, there is an automorphism $\alpha$ of $A$ that is homotopic to the identity and satisfies
@@
This in particular implies that the ‘noncommutative’ transport constant $k_{C(X)}$ of $(C(X),\rho)$ (in the sense of Definition 6.1, relative to any nucleus $\mathcal{L}(C(X))$) agrees with the ‘commutative’ transport constant $k_{X}$ introduced in [Jac23b, Definition 2.6]. Actually, that is not quite true: Definition 6.1 is stated for faithful measures while [Jac23b, Definition 2.6] only considers measures that are good, that is, faithful as well as diffuse. So, formally, we have $k_{C(X)}\geq k_{X}$. The difference really becomes substantive when one is interested in an exact transport map rather than an approximate one, and for our present needs is insignificant. Indeed, as observed in [Jac23b, Remark 2.7], our typical method of transport does not depend on diffuseness (see the proof of Theorem 6.4), so there is no practical difference between the two notions. Moreover, as good measures are generic in $\operatorname{\mathcal{M}}(X)$, faithful measures can be perturbed to good ones (see the proof of [Jac23b, Theorem 3.1]). A final point in favour of working with faithful traces, as discussed at the beginning of this section, is that these are what actually parameterise embeddings.
@@
Let $X$ be a compact, connected Riemannian manifold of dimension at least three and let $\rho$ be an intrinsic metric on $X$ that is Lipschitz equivalent to the intrinsic metric induced by the Riemannian structure. Let $X_{p,q}$ be a generalised dimension drop algebra over $X$ for some $p,q\in\mathbb{N}$ and $x_{0}\neq x_{1}\in X$. Then, $k_{(X_{p,q},\rho)}=1$ relative to $\mathcal{D}_{r}(X_{p,q})$.
@@ 6.2. Distances between unitary orbits
$d_{U}(\varphi,\psi)|_{\mathcal{L}(A)}:=\adjustlimits{\sup}_{F\subset\joinrel\subset\mathcal{L}(A)}{\inf}_{u\in\mathcal{U}(B)}\sup_{a\in F}\|\varphi(a)-u\psi(a)u^{*}\|.$ (6.10)
@@
The inequality (6.12) is proved in exactly the same way as [JSV21, Theorem 4.11] and its generalisation [Jac23b, Theorem 3.1], which cover commutative domains $A=C(X)$. (Briefly, classification [CGSTW23, Theorem B] lets us convert transport maps $\alpha\colon A\to A$ into conjugating unitaries $u\in B$.) For (6.13), since we know from Theorem 6.4 that $k_{(A,\rho)}=1$ relative to the canonical nucleus $\mathcal{D}_{r}(A)$ defined in (6.4), it then suffices to prove that $d_{U}(\varphi,\psi)|_{\mathcal{D}_{r}(A)}\geq W_{\infty}(\varphi,\psi)$. This can also be deduced from [JSV21, Jac23b], since the embedding $\iota$ of $C(X)=X_{1,1}$ into the centre of $A=X_{p,q}$ restricts to an inclusion of canonical nuclei. It follows that
@@
[…]
@@
$\overrightarrow{W}_{\infty,\rho}((\sigma_{k})_{k=1}^{\infty},(\tau_{k})_{k=1}^{\infty})=\limsup_{k\to\infty}\overrightarrow{W}_{\infty,\rho_{k}}(\sigma_{k},\tau_{k})=\sup_{k\in\mathbb{N}}\overrightarrow{W}_{\infty,\rho_{k}}(\sigma_{k},\tau_{k}).$ (6.18)
@@ 6.3.2. Empirical embeddings
Funding
Competing interests
The author has none to declare.